# Predicting the structure of dynamic graphs

### 08 Jan 2024 v1

@@ Abstract
Dynamic graph embeddings, inductive and incremental learning facilitate predictive tasks such as node classification and link prediction. However, predicting the structure of a graph at a future time step from a time series of graphs, allowing for new nodes has not gained much attention. In this paper, we present such an approach. We use time series methods to predict the node degree at future time points and combine it with flux balance analysis – a linear programming method used in biochemistry – to obtain the structure of future graphs. Furthermore, we explore the predictive graph distribution for different parameter values. We evaluate this method using synthetic and real datasets and demonstrate its utility and applicability.
@@ 1 Introduction
Networks are everywhere; from protein-protein interaction networks in biology to social networks on social media platforms, we are surrounded by networks in different mediums and scales. As such, there is a extensive body of research on different aspects of networks/graphs, their embeddings and prediction. Our focus is on predicting the structure of a dynamic, undirected, unweighted network/graph at a future time step including new nodes and edges.
Dynamic graphs such as those encountered in social networks change over time with the addition and deletion of new nodes and edges. In contrast, the structure of static graphs remain constant. For example, while the traffic changes over time in road networks, the underlying graph structure remains the same. Even though static graphs have been studied for longer, there is a lot of interest on dynamic graphs (Pareja et al., 2020; Trivedi et al., 2019; You et al., 2022; Barros et al., 2021; Sankar et al., 2020). The broad applicability of dynamic graphs to many scenarios including recommender systems, autonomous vehicles and social networks make it an expanding research field.
@@
Generally dynamic graphs are represented in two different formats: discrete-time dynamic graphs (DTDG) and continuous-time dynamic graphs (CTDG) (Kazemi et al., 2020). Discrete-time dynamic graphs consider a sequence of graphs at regular time intervals, for example, graphs on social network daily activity. Continuous-time dynamic graphs (CTDG) are more event-driven, where each event can be the formation of an edge, addition of a new node or a similar activity. We focus on discrete-time dynamic graphs (DTDG) in this work.
One of the most popular tasks considered in dynamic graphs is node classification. Predicting the paper subject categories or classifying protein roles in a protein-protein interaction network are node classification examples. Other tasks include graph classification, link prediction, edge classification, entity/relation prediction, time prediction and network clustering. Most tasks consider a sequence of graphs for $i\in\{1,\ldots,t\}$ and predict the required property of the graph at $t+1$ using some information about the graph at $t+1$. For node classification, input features of nodes at time $t+1$ are generally considered when predicting the category of the nodes (Pareja et al., 2020). Link prediction considers if a link exists between two nodes $u$ and $v$. Thus, the nodes at time $t+1$ are known. Dynamic entity/relation prediction predicts a missing entry in the tuple $(u,r,v)$ , where $u$ and $v$ are nodes and $r$ is the relation (Kazemi et al., 2020). In all these scenarios, some information about the graph at time $t+1$ is known. Our work is different from these tasks because we do not know anything about the graph at time $t+1$. From a sequence of graphs, we predict the structure – nodes and edges – of the graph at a future time step. To the best of our knowledge, this task has not been studied before. In this study we consider simple graphs, i.e., unweighted, undirected graphs without any node or edge attributes.
@@
The methods we use are motivated from Flux Balance Analysis (FBA) (Orth et al., 2010) and combined with time series prediction. FBA is a mathematical approach widely used in biochemistry to reconstruct metabolic networks by using linear and mixed-integer programming. By solving an optimization problem maximizing the biomass function subject to a large number of constraints, FBA arrives at the flux solution describing a network of chemical reactions. We adapt FBA to a graph/network prediction framework by considering the degree of each node instead of the flux of chemical reactions. To predict the graph at a future time step, we incorporate the predicted degree of each node in the constraints. Thus, the adapted FBA solution gives us a graph satisfying the degree constraints at a future time step; it predicts the structure of the future graph.
2 Related Work
2.1 Inductive graph learning
Inductive graph learning focuses on changing graph structure and the ability to generalize to unseen nodes. In an inductive setting, all nodes may not be present at the training stage. This is different to transductive learning, where the graph structure used in training is used for testing. Existing work on transductive graph learning and embedding approaches far exceeds that of inductive learning. Inductive learning is mostly used for node classification (Hamilton et al., 2017; Su et al., 2023; Bai et al., 2019; Rossi et al., 2020; Yang et al., 2016). Yang et al. (2016) explain the difference at an architectural level considering the inputs into account. The transductive formulation takes the input features as well as the graph embedding as inputs to predict the class label. Their proposed inductive algorithm Planetoid-I takes only the input features and defines the embedding as a prametrizatized function of the features, which then is used to predict the class label. Citation networks are a popular application of inductive graph learning, where new paper subject categories are predicted using features but without the knowledge of the new graph. Other applications include similarity ranking and graph visualization (Bai et al., 2019) and within-network link classification and link prediction (Rossi et al., 2020) where a graph with a fraction of missing edges is used.
2.2 Incremental and dynamic graph learning
In dynamic network settings such as social networks, the graphs are large and it is computationally prohibitive to retrain a full model at each time step. Thus, it is important to update the model without retraining it from scratch. Due to the significance of the research problem and its strong pertinence, both incremental and dynamic graph learning have attracted a lot of attention (Kim et al., 2022; Tan et al., 2022; Xu et al., 2020; Yin et al., 2021; Yuan et al., 2021; Zhang et al., 2022; Galke et al., 2021; Wang et al., 2020; Han, Karunasekera & Leckie, 2021).
Both Incremental and dynamic graph learning encompass a set of methods that learn from a sequence of graphs and generalize to unseen graphs. However, the two methodologies focus on different aspects of learning. Incremental graph learning also known as continual or lifelong graph learning focuses on overcoming the catastrophic forgetting problem – the significant decrease of model performance of previous tasks after being trained on new tasks (Xu et al., 2020). Methods employed to combat catastrophic forgetting in incremental graph learning can be categorized into three types; regularization-based methods, memory-replay based methods, and parametric isolation based methods. Each category of methods strive to resolve catastrophic forgetting in different ways, either by penalizing changes to parameter updates, selecting representative data from previous tasks or by introducing new parameters to avoid drastic changes.
Dynamic graph learning has a different emphasis. The focus of dynamic graph learning is to capture the temporal dynamics and keep the graph representations up to date (Zhang et al., 2022; Su et al., 2023). The catastrophic forgetting problem is not the focus of dynamic graph learning. While these two topics have different technical emphasis, both incremental and dynamic graph learning train models to perform similar tasks. Two popular tasks for both scenarios are node classification and graph classification. In both settings new nodes can appear and they can belong to new classes as well as existing classes. Similarly new categories of graphs can appear over time.
2.3 Link prediction in dynamic graphs
Link prediction in graphs has a long history (Getoor et al., 2003). The methods used for link prediction can be broadly categorized into probabilistic, spectral clustering, time series, deep learning including embedding and graph neural networks, matrix factorization and other methods (Divakaran & Mohan, 2020). Recent work includes link prediction using neural ordinary differential equations (Luo et al., 2023; Han, Ding, Ma, Gu & Tresp, 2021).
Most work done on link prediction consider the set of nodes as fixed and known beforehand, while the links are unknown (Chen et al., 2022, 2021; Qu et al., 2020; Ma et al., 2018). In this setting, if a node $v_{i}$ appears at time $t_{j}$ and makes its first link with node $v_{k}$, the modeling paradigm assumes the existence of the node from time $t_{0}<t_{j}$ and treats $v_{i}$ as a silent node that makes its first link at time $t_{j}$. In the other setting the set of nodes is more dynamic and not completely known at the beginning. In this setting, they consider the problem “given these nodes, are there links between them?” and evaluate performance on a hidden set of links (Goyal et al., 2017). Thus, the dynamic link prediction problem predicts the links given the nodes at a specific time.
While link prediction, inductive, incremental and dynamic graph learning are related, our work is different as we predict the graph structure including new nodes and edges at a future time point.
@@ 3 Methodology
3.1 Notation
We consider a series of undirected, unweighted, growing graphs $\mathcal{G}_{t}=\left(V_{t},E_{t}\right)$ without loops, where $V_{t}\subseteq V_{t+1}$ (also known as projective (Cai et al., 2016)) for time $t\in\{1,\ldots,T\}$ and focus on predicting $\mathcal{G}_{T+h}=\left(V_{T+h},E_{T+h}\right)$. Let $V_{t}=\{v_{1},\ldots,v_{n_{t}}\}$ denote the set of vertices in graph $\mathcal{G}_{t}$ with $|V_{t}|=n_{t}$. Let $e_{ij,t}$ denote the edge between nodes $i$ and $j$ at time $t$, where $i$ and $j$ can denote new or old vertices. Thus, the set of edges in $\mathcal{G}_{t}$ is $E_{t}=\{e_{ij,t}\}_{i,j\in V_{t}}$. Let $d_{i,t}$ denote the degree of vertex $v_{i}$ in $\mathcal{G}_{t}$, i.e., in a unweighted graph setting $d_{i,t}=\sum_{j}e_{ij,t}$. We denote all predictions by a hat $\hat{}$. For example, the predicted number of vertices at time $T+h$ is denoted by $\hat{n}_{T+h}$.
Our aim is to predict $\mathcal{G}_{T+h}=\left(V_{T+h},E_{T+h}\right)$, i.e., the structure of the graph at $T+h$ including new vertices and new edges. In particular we describe, in some form, the distribution of graphs at $t=T+h$ which can include a variable number of nodes and edges. Our approach to predicting $\mathcal{G}_{T+h}$ comprises two steps: 1. Predict the number of vertices and their degree distributions at $T+h$ and 2. predict the edges using the predicted degree distributions taking into account the new vertices.
The degree of each vertex $d_{i,t}$ in $\mathcal{G}_{t}$ can be considered as a time series $\{d_{i,t}\}_{t=t_{0,i}}^{T}$ starting from time $t_{0,i}$ when the $i$th vertex appeared. Similarly the time series $\{n_{t}\}_{t=1}^{T}$ keeps track of the number of vertices in $\mathcal{G}_{t}$. We can predict $d_{i,T+h}$ and $n_{T+h}$ using standard time series methods. Let us call our predictions $\hat{d}_{i,T+h}$ and $\hat{n}_{T+h}$ respectively. If $\hat{n}_{T+h}>n_{T}$ we can add $\hat{n}_{T+h}-n_{T}$ new nodes to the graph. From the time series of graphs $\left\{\mathcal{G}_{t}\right\}_{t=1}^{T}$ we know the new nodes at each time step and their degree distribution. We take the average degree of new nodes as the predicted degree $\hat{d}_{i,T+h}$ for new nodes $v_{i}$. Thus, we have a degree prediction for all nodes, old and new. Then the main question is how to determine the structure of the graph in terms of edges taking into account the predicted degree distribution. Distributing the edges among the vertices satisfying the degree distribution is an optimization problem. We use Flux Balance Analysis (Orth et al., 2010), a methodology that uses linear and integer programming to reconstruct metabolic networks, to allocate the edges and thus obtain the predicted network.
3.2 Flux Balance Analysis: a quick introduction
Flux Balance Analysis (FBA) is a mathematical approach widely used in biochemistry to reconstruct metabolic networks from partial information. A large pool of chemical reactions is used as input in this process. By mathematically representing chemical reactions using stoichiometric coefficients, FBA comes up with a set of constraints that need to be satisfied. The solution space defined by the set of constraints describing the potential network is denoted by the matrix equation $\bm{S}\bm{v}=0$ along with the inequalities $a_{i}\leq v_{i}\leq b_{i}$ for $i\in\{1,\ldots,q\}$ (Orth et al., 2010), where $\bm{v}$ denotes the vector of fluxes to be determined and $\bm{S}$ denotes the constraint matrix. As $\bm{S}$ is a $p\times q$ matrix with $p<q$, there are multiple solutions to this system of linear equations. FBA selects the optimal solution by maximizing a growth function denoted by $Z=\sum_{i}c_{i}v_{i}$. We note that FBA is one instance where linear programming is used in a graph/network setting. In fact, linear programming is used in many network flow applications including traffic modeling in road and communication networks and oil flow modeling in pipeline networks (Bazaraa et al., 2011).
3.3 Adapting linear and integer programming to graph prediction
@@
In our scenario we are not modeling a flow. However, we can use linear and integer programming to solve the problem of graph prediction. We consider an optimization problem where the constraints describe the degree of vertices and each variable denotes the presence/absence of an edge. The sum of the edges of a vertex equals its degree. But as we are predicting the graph at time $T+h$ we do not know the degree of the vertices. This is not a problem because we can use the prediction $\hat{d}_{i,t}$ instead. Using the prediction $\hat{d}_{i,t}$ forces us to consider the type of constraints we use, i.e., if we employ strict equalities with each equation specifying the degree of a vertex, then it is possible that such a graph cannot exist because the constraints force the solution space to be empty. For example an undirected, unweighted graph without loops of 3 vertices with degrees 2, 2, 1 cannot exist, but with degrees 2, 1, 1 or 2, 2, 2 can exist. Therefore, we need to incorporate some variability that enables the existence of feasible graphs. This can be achieved by using inequalities instead of strict equalities and by specifying reasonable upper bounds for the vertex degree, such as the upper confidence bound of the predicted degree. These adjustments can ensure a non-empty solution space. Thus, a constraint is of the form
$\sum_{j}\hat{e}_{ij,T+h}\leq f_{u}\left(\hat{d}_{i,T+h}\right)$ (1)
where $f_{u}$ denotes an upper bound function. To denote these constraints using matrix notation we introduce the matrix $S$, which we obtain from a graph’s adjacency matrix as follows:

### 25 Jul 2024 v2

# Forecasting the structure of dynamic graphs

@@ Abstract
Many aspects of graphs have been studied in depth. However, forecasting the structure of a graph at future time steps incorporating unseen, new nodes and edges has not gained much attention. In this paper, we present such an approach. Using a time series of graphs, we forecast graphs at future time steps. We use time series forecasting methods to predict the node degree at future time points and combine these forecasts with flux balance analysis – a linear programming method used in biochemistry – to obtain the structure of future graphs. We evaluate this approach using synthetic and real-world datasets and demonstrate its utility and applicability.
@@ 1 Introduction
We are surrounded by networks in different mediums and scales. From a biological perspective protein-protein interaction networks mediate cellular responses much needed for our biological function; from an infrastructure viewpoint, road, rail and airline networks govern a significant portion of our time spent on commuting, and from an online aspect, social media networks have been part of our lives since the latter part of the twentieth century. There is extensive research conducted on different aspects of networks/graphs including link prediction, anomaly detection, time series forecasting, inference and embeddings to name a few. In this paper, we focus on forecasting the structure of dynamic, undirected, unweighted graphs at future time steps including new nodes and edges.
Dynamic graphs, such as those encountered in social networks change over time with the addition and deletion of nodes and edges. When people join a social network new nodes are added, and when they leave nodes should ideally be deleted. Similarly, networks grow in epidemiological settings, where human movement and proximity play a big role. A somewhat different example is a home electricity network, where different appliances can plug in to different power sockets at different times for different durations. If each type of appliance is modelled as a separate class of nodes then we can model the network as a heterogenous graph. Nodes such as phones move over the network when appliances connect to different power outlets at different times. Going beyond the home electricity network, this pattern is seen when electric cars charge at different locations at different time points. These are examples of dynamic graph structures. In contrast, the structure of static graphs remains constant. For example, while the traffic changes over time in road networks, the underlying graph structure remains the same apart from incidents of road closures. While static graphs have been studied for longer, there is a lot of interest in dynamic graphs.
@@
Generally dynamic graphs are represented in two different formats: discrete-time dynamic graphs (DTDG) and continuous-time dynamic graphs (CTDG) (Kazemi et al., 2020). Discrete-time dynamic graphs consider a sequence of graphs at regular time intervals, for example, daily graphs on social network activity. Thus, the data comes in the form of a time series of graphs. In contrast, continuous-time dynamic graphs (CTDG) are more event-driven, where each event can be the formation of an edge, the addition of a new node or similar activity. Timestamped edgelists are generally used in the continuous setting. Our focus on discrete-time dynamic graphs (DTDG) and as such we work with time series of graphs.
Of the diverse methodologies used for graph prediction tasks, our work has some parallels with network time series forecasting, which considers time series of node attributes and forecasts them taking into account the network structure. In this context, a node attribute $x_{i,t}$ is observed for each node $i$ in the network for all time stamps $t$ and the task is to forecast the attribute at future time points (Kim & Oh, 2024). In network time series forecasting, the structure of the network is fixed and known. In our case, we forecast the structure of the network at a future time point. To the best of our knowledge, this task has not been studied before. In addition to network time series forecasting, network structure forecasting has some similarities with link prediction. However, our work is different because link prediction considers if a link exists between two nodes $u$ and $v$ where both nodes $u$ and $v$ are known to exist in the future network. Thus, the nodes in the network at time $t+1$ are known in link prediction. Furthermore, link prediction studies divide links into training and test sets and trains a model on the links in the training set and tests it on the unseen links (Kumar et al., 2020). Link prediction does not forecast the structure of the full network, which is our focus.
Forecasting the structure of dynamic networks have applications in energy, social and disease networks. While traditionally energy networks focused on aggregate measurements, recent advances such as smart energy initiatives need more adaptable techniques and forecasts. For example, behind the meter forecasting and monitoring (Erdener et al., 2022; Bauknecht et al., 2020) and peer-to-peer trading in energy networks (Alam et al., 2019) call for newer methods as the network can constantly evolve. Smart homes can act as energy providers coining the term “prosumers” in energy network literature. Not only electric vehicles move locations, but they can act as power sources changing the network structure (Alshahrani et al., 2019). In social networks, the network structure is dynamic and network growth and evolution are topics of interest (Sobkowicz et al., 2012; Wu et al., 2017) especially as they can model responses to certain events. Similarly, modeling disease spread in a network by contact (Read & Keeling, 2003) and the evolution of infectious diseases (Leskovec et al., 2007) consider dynamic network structures.
@@
In this paper, we consider simple graphs, i.e., unweighted, undirected graphs without any node or edge attributes. The proposed method is motivated by Flux Balance Analysis (FBA) (Orth et al., 2010). FBA is a mathematical approach widely used in biochemistry to reconstruct metabolic networks by using linear and mixed-integer programming. By solving an optimization problem maximizing the biomass function subject to a large number of constraints, FBA arrives at the flux solution describing a network of chemical reactions. We adapt FBA to a graph/network forecasting framework by considering the degree of each node instead of the flux of chemical reactions. To forecast the graph at a future time step, we incorporate the degree forecast of each node in the constraints. Thus, the adapted FBA solution gives us a graph satisfying the degree forecasts; this is the graph forecast. Our contributions are summarized as follows:
  • We address the challenge of forecasting the evolving structure of dynamic graphs accounting for the emergence of new nodes and edges at future time steps.
  • We propose a novel approach that adapts flux balance analysis (a method used to reconstruct metabolic networks in biochemistry) enabling graph forecasting.
  • We demonstrate the applicability of our method on both synthetic and real-world datasets.
2 Related Work and Notation
In this section we briefly introduce Flux Balance Analysis (FBA) and introduce the notation.
2.1 Flux Balance Analysis: a quick introduction
Flux Balance Analysis (FBA) is a mathematical approach widely used in biochemistry to reconstruct metabolic networks from partial information. A large pool of chemical reactions is used as input to this approach. By mathematically representing chemical reactions using stoichiometric coefficients, FBA comes up with a set of constraints that need to be satisfied. The solution space defined by the set of constraints describing the potential network is denoted by the matrix equation $\bm{S}\bm{v}=0$ along with the inequalities $a_{i}\leq v_{i}\leq b_{i}$ for $i\in\{1,\ldots,q\}$ (Orth et al., 2010), where $\bm{v}$ denotes the vector of fluxes to be determined and $\bm{S}$ denotes the constraint matrix. As $\bm{S}$ is a $p\times q$ matrix with $p<q$, there are multiple solutions to this system of linear equations. FBA selects the optimal solution by maximizing a growth function denoted by $Z=\sum_{i}c_{i}v_{i}$.
2.2 Notation
We consider a time series of undirected, unweighted, graphs without loops $\mathcal{G}_{t}=\left(V_{t},E_{t}\right)$ where $V_{t}$ and $E_{t}$ denote the set of vertices and edges respectively. We consider the scenario of growing graphs, where $V_{t}\subseteq V_{t+1}$ and focus on forecasting $\mathcal{G}_{T+h}=\left(V_{T+h},E_{T+h}\right)$. Let $V_{t}=\{v_{1},\ldots,v_{n_{t}}\}$ denote the set of vertices in graph $\mathcal{G}_{t}$ with $|V_{t}|=n_{t}$. Let $e_{ij,t}$ denote the edge between nodes $i$ and $j$ at time $t$, where $i$ and $j$ can denote new or old vertices and let $m_{t}=|E_{t}|$ denote the number of edges in $\mathcal{G}_{t}$. Let $d_{i,t}$ denote the degree of vertex $v_{i}$ in $\mathcal{G}_{t}$, i.e., in an unweighted graph setting $d_{i,t}=\sum_{j}e_{ij,t}$. The edge-density of graph $\mathcal{G}_{t}$ is given by $2m_{t}/(n_{t}(n_{t}-1))$ which we denote by $\text{dens}(t)$.
We denote the union of graphs by $\mathcal{G}_{U}=\mathcal{G}_{1}\bigcup\ldots\bigcup\mathcal{G}_{T}$ and the set of $K$ most popular nodes (largest degree) in $\mathcal{G}_{U}$ by $\kappa$. When computing $\mathcal{G}_{U}$ multiple edges between the same vertices are collapsed to a single edge, i.e., we do not consider multiple edge graphs. With a slight abuse of notation we denote the degree of node $i$ in $\mathcal{G}_{U}$ by $d_{i,U}$. We denote all forecasts by a hat $\hat{}$. For example, the forecast number of vertices at time $T+h$ is denoted by $\hat{n}_{T+h}$. Our aim is to obtain a forecast for $\mathcal{G}_{T+h}=\left(V_{T+h},E_{T+h}\right)$, which we denote by $\hat{\mathcal{G}}_{T+h}=\left(\hat{V}_{T+h},\hat{E}_{T+h}\right)$ containing the structure of the graph at $T+h$ including new vertices and new edges.
@@ 3 Methodology
FBA is generally used to model the flux, which can be thought of as a flow. In our scenario, we are not modelling a flow. As such, we adapt FBA to solve the problem of graph forecasting. Figure 1 gives a flowchart of the proposed methodology.
3.1 Adapting FBA to graph forecasting
@@
In metabolic networks FBA is used for network reconstruction, not forecasting. We adapt FBA for graph forecasting by using degree forecasts of individual nodes. We consider an optimization problem where the constraints describe the degree of vertices and each variable denotes the presence/absence of an edge. The sum of the edges of a vertex equals its degree. But as we are forecasting the graph at time $T+h$ we do not know the degree of the vertices at time $T+h$. This is not a problem because we can use the forecasts $\hat{d}_{i,T+h}$ instead. Using the forecasts $\hat{d}_{i,T+h}$ forces us to consider the type of constraints we use, i.e., if we employ strict equalities with each equation specifying the degree of a vertex, then it is possible that such a graph cannot exist because the constraints force the solution space to be empty. For example, an undirected, unweighted graph without loops of 3 vertices with degrees 2, 2, 1 cannot exist, but with degrees 2, 1, 1 or 2, 2, 2 can exist. Therefore, we need to incorporate some variability that enables the existence of feasible graphs. This can be achieved by using inequalities instead of strict equalities and by specifying reasonable upper bounds for the vertex degree, such as the upper confidence bound of the degree forecast. These adjustments can ensure a non-empty solution space. Thus, a constraint is of the form,
$\sum_{j}\hat{e}_{ij,T+h}\leq f_{u}\left(\hat{d}_{i,T+h}\right),$ (1)
where $f_{u}$ denotes an upper bound function. Typically we use the upper confidence limit of the degree forecast as $f_{u}$.
Figure 1: The graph time series $\{\mathcal{G}_{t}\}_{t=1}^{T}$ is used to extract multiple time series: $n_{t}$ is the number of nodes in $\mathcal{G}_{t}$ and $d_{i,t}$ is the degree of each node $i$ at time $t$. The forecast $\hat{n}_{T+h}$ is computed using the time series $\{n_{t}\}_{t=1}^{T}$ and the forecast of the degree of existing nodes $\hat{d}_{i,T+h}$ is computed using the time series $\{d_{i,t}\}_{t=1}^{T}$ using ARIMA models. The mean degree of nodes that appear for the first time, $d_{avg}$ is computed. The number of new nodes $n_{new}=\hat{n}_{T+h}-n_{T}$. Using $\{\mathcal{G}_{t}\}_{t=1}^{T}$ the union graph $\mathcal{G}_{U}$ and matrix $S$ are computed. Coefficients for optimization $\xi_{ij}$ are determined depending on the scheme. Finally, adapted Flux Balance Analysis is carried out giving $\hat{\mathcal{G}}_{T+h}$ as the output.
3.2 Time series extraction and forecasting
In our adaptation, FBA uses time series forecasts, the matrix $S$ and coefficients $\xi_{ij}$ as inputs. The first step is to extract certain time series from $\{\mathcal{G}_{t}\}_{t=1}^{T}$ and forecast them at $T+h$. The degree of each vertex $d_{i,t}$ in $\mathcal{G}_{t}$ can be considered as a time series $\{d_{i,t}\}_{t=t_{0,i}}^{T}$ starting from time $t_{0,i}$ when the $i$th vertex appears. Similarly the time series $\{n_{t}\}_{t=1}^{T}$ keeps track of the number of vertices in $\mathcal{G}_{t}$. We can forecast each time series $\{d_{i,t}\}_{t=t_{0,i}}^{T}$ and obtain a point forecast and its confidence intervals using standard time series methods. We use ARIMA models with automatic parameter selection (Hyndman & Khandakar, 2008) to obtain the point forecasts and their upper confidence limits. Let us call our point forecasts $\hat{d}_{i,T+h}$ and $\hat{n}_{T+h}$ respectively. Let $u$ denote the percentile used to obtain the upper confidence bound of $\hat{d}_{i,T+h}$, which we call $f_{u}(\hat{d}_{i,T+h})$. While we can use a percentile $\gamma$ to obtain an upper bound of $\hat{n}_{T+h}$, we generally use the point forecast $\hat{n}_{T+h}$ corresponding to $\gamma=0.5$. If $\hat{n}_{T+h}>n_{T}$ we can add $\hat{n}_{new}=\hat{n}_{T+h}-n_{T}$ new nodes to the graph.
From the time series of graphs $\left\{\mathcal{G}_{t}\right\}_{t=1}^{T}$ we know the new nodes at each time step and their degree distribution. We take the average degree of new nodes $d_{avg}$ as the forecast degree $\hat{d}_{i,T+h}$ for new, unseen nodes $v_{i}$. Thus, we have degree forecast for all nodes, old and new.
3.3 The matrix $S$ and the union graph $\mathcal{G}_{U}$
For a graph with $n$ vertices, we consider a zero-inflated incidence matrix $S$ of size $n\times n(n-1)/2$. The incidence matrix of a graph has $n$ rows and $m$ columns where $n$ denotes the number of nodes and $m$ the number of edges. The $ij$th entry of the matrix is set to $1$ if vertex $v_{i}$ is incident with edge $e_{j}$. Our version of the incidence matrix is inflated with zero-columns to incorporate edges that are not present in the graph. We obtain $S$ from the adjacency matrix $A_{n\times n}$ as follows: The columns represent the edges obtained from unrolling the upper triangular part of $A$ row by row, and the rows of $S$ represent the vertices 1 to $n$. If an edge $e_{ij}$ is present, i.e., $e_{ij}=1$ with $i<j$ then $s_{ik}=s_{jk}=1$ where $k$ denotes the column corresponding to edge $e_{ij}$, with $k=\left((n-1)+(n-2)+\ldots+(n-i+1)\right)+(j-i)=\frac{1}{2}(i-1)(2n-i)+(j-i)$.
The matrix $S$ for forecasting the graph at $T+h$ is obtained by considering the union graph from the time series of graphs and new vertices as per the forecast $\hat{n}_{T+h}$.